# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v1

@@ 4.1 Setup
where $R_{n}=\text{diag}(0,0.3,0.3,\ldots,0.3)$. Under this $R_{n}$ the error term $u_{t}$ in the main regression (2.1) remains i.i.d. satisfying the m.d.s. condition in Assumption 1, while the unit root generator $\Delta X_{t}$ and the stationary regressors $Z_{t}$ are AR(1) processes. The covariance matrix $\Sigma=(\Sigma_{ij})_{i,j\in[p]}$ is specified as
@@
Recall $0_{p}$ is a $p$-dimensional zero vector. The true coefficient vectors are:
@@
Both XDlasso and Dlasso involve Slasso, where the selection of tuning parameters $\lambda$ and $\mu$ affects finite sample performance. In our experiments, we employ the block 10-fold cross-validation (CV), splitting the sample into 10 equally sized chronologically ordered consecutive blocks for validations. In Theorem 1, the tuning parameters are specified as constants multiplied by the appropriate rate of convergence determined by the sample size $n$, dimensionality $p$ and the mixing condition constant $r$. As a benchmark, we also calibrate the tuning parameters following Lee et al., (2022) to examine the validity of the theoretical orders of tuning parameters specified in Theorem 1. Specifically, we perform $500$ pilot replications for each DGP, with $n_{0}=400$ , $\left(p_{x0},p_{z0}\right)=\left(100,150\right)$, and $p_{0}=p_{x0}+p_{z0}$. In each replication $q=1,2,\dots,500$, we use the 10-fold cross-validation to choose the tuning parameters $\lambda^{(q)}$ and $\mu^{(q)}$, and calibrate the constants as
$C_{\lambda}^{\left(q\right)}=\lambda^{(q)}n_{0}^{1/2}/(\log p_{0})^{\frac{3}{2}+\frac{1}{2r}},\ \ C_{\mu}^{\left(q\right)}=\mu^{(q)}n_{0}^{[(1-\tau)\wedge\tau]/2}/(\log p_{0})^{2+\frac{1}{2r}},$
where $r=1$ and $\tau=0.5$ are chosen in the simulation. We then fix $\widehat{C}_{\star}=\mathrm{median}(C_{\star}^{\left(1\right)},\ldots,C_{\star}^{\left(500\right)})$ for $\star\in\left\{\lambda,\mu\right\}$ in the full-scale experiments. The tuning parameters are then set as $\widehat{\lambda}=\widehat{C}_{\lambda}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\widehat{\mu}=\widehat{C}_{\mu}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ as in Theorem 1.
@@ 5 Empirical Applications
In recent years, high dimensional macroeconomic data have been extensively used to forecast key variables of interest; see Smeekes and Wijler, (2018), Medeiros et al., (2021), and Giannone et al., (2021), for examples. Researchers have primarily focused on the point estimation of forecast. There has been limited empirical exploration of statistical inference on predictive power of specific predictors, due to a lack of suitable toolkits. This section showcases two empirical applications using our proposed XDlasso inference method. We utilize the monthly data of 112 U.S. macroeconomic variables spanning from January 1960 to April 2024, sourced from the FRED-MD dataset by McCracken and Ng, (2016).
@@ 5.1 Predictability of Stock Return Using Earnings-Price Ratio
In spite of a prominent economist and statistician with a national reputation, Irving Fisher suffered devastating personal finance losses during the 1929 Great Depression (Dimand and Geanakoplos, , 2005). Were he alive today, he would have mixed feelings about the fact that return predictability remains inconclusive after a century of extensively studies. Among the various predictors examined, the predictive power of valuation ratios, particularly the earnings-price ratio, has been subject to widespread discussion. In this section, we investigate the predictability of stock returns using the log earnings-price ratio while incorporating high dimensional mixed-root control covariates.
@@ 5.1.1 Data
Figure B.1 illustrates the time series plot of the monthly return of the S&P 500 index and the log earnings-price ratio from January 1960 to April 2024. The log earnings-price ratio exhibits persistent patterns, while the S&P 500 return appears stationary. We further report the AR(1) coefficient estimates and Augmented Dickey–Fuller (ADF) test $p$-values of both series under different sample periods in Table B.1. The S&P 500 return is evidenced to be stationary under the full sample with an ADF test $p$-value below 1%, rejecting the null hypothesis of nonstationarity. Conversely, the log earnings-price ratio shows high persistence with an AR(1) coefficient estimate equal to $0.993$. Given the $p$-value of $0.081$, nonstationarity is not rejected under the 5% significance level.
@@
Our analysis incorporates high dimensional covariates including all other 110 macroeconomic variables from the FRED-MD dataset. These variables are mixed with stationary and nonstationary time series. To facilitate the stationarization, FRED assigns a transformation code (TCODE), as recommended transformation of potentially nonstationary time series into stationary ones. Given the stationary pattern of the dependent variable, in our analysis, we follow the TCODE to transform the 110 time series and subsequently use the transformed variables as covariates. Nevertheless, we highlight that these elementary transformations are not a silver bullet in getting rid of nonstationarity. We perform the ADF test for each of the transformed variables for different sample periods. A nontrivial proportion (9.1%) of the transformed variables still demonstrate nonstationarity based on the ADF test at 5% significance level for both Pre-1994 and Post 1994 subperiods. The high persistence of the log earnings-price ratio and other covariates suggests the necessity of XDlasso.
@@ 5.1.2 Results
We study the one-month-ahead regression $\text{Return}_{t}=\alpha+\theta_{1}^{*}\times\text{logEP}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes a high dimensional vector that collects all control variables. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the log earnings-price ratio in forecasting the S&P 500 monthly return. In addition to XDlasso and Dlasso, we conduct the IVX inference in a simple regression setup, omitting all control covariates, to demonstrate the necessity of incorporating high dimensional control covariates in practice. We compare the results of XDlasso to those of Dlasso for the high dimensional regression and IVX for a simple regression $\text{Return}_{t}=\alpha+\theta_{1}^{*}\times\text{logEP}_{t-1}+u_{t}.$
@@ 5.2 Predictability of Inflation Using Unemployment Rate
It is essential for monetary policymakers to understand the relationship between unemployment and inflation. As Engemann, (2020) pointed out, “The Federal Reserve has a dual mandate to promote maximum sustainable employment and price stability.” First alluded by Fisher, (1926, 1973) though, Phillips, (1958) popularized the Phillips curve — a negative relationship between the level of unemployment and the change rate of money wage rates. There has been a prolonged debate about whether the unemployment is a credible barometer for inflation. We revisit the Phillips curve in a predictive regression framework utilizing the FRED-MD dataset.
@@ 5.2.1 Data
Similar to our first empirical application in Section 5.1, we follow Benati, (2015) to delineate three subperiods in addition to the full sample: Pre-Volcker (January 1960–July 1979), Volcker and Greenspan (August 1979–January 2006), and Bernanke, Yellen, and Powell (February 2006 - April 2024). These subperiods correspond to different eras in U.S. monetary policy, each named after the Federal Reserve chairperson who presided during that time. This periodization allows us to examine how the relationship between inflation and unemployment may have evolved across different policy regimes and economic conditions.
@@ 5.2.2 Results
We study, in our predictive regression framework, a one-month-ahead regression $\text{Inflation}_{t}=\alpha+\theta_{1}^{*}\times\text{Unrate}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes all transformed covariates. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the unemployment rate in forecasting inflation. The predictive form is of particular interest among policymakers in leveraging the relationship as a practical tool.
@@
Table 4 reports the point estimates and standard errors for $\theta_{1}^{*}$ using IVX, Dlasso and XDlasso across sample periods. The results by XDlasso in different subperiods are unanimous that the unemployment rate does not have significant predictive power in forecasting the inflation rate across all periods. These results add new insight to the recent debates on the Phillips curve: the predictive power of unemployment for inflation appears to be diminished when we control for other high dimensional time series as covariates. In contrast, IVX and Dlasso yield contradictory results. Dlasso delivers significantly positive estimates for the full sample and the Pre-Volcker period at the 1% significance level. Conversely, IVX detects significant negative effects for the full sample and Bernanke-Yellen-Powell period at the 5% and 10% significance levels, respectively. These discrepancies may be attributed to the failure to control for confounders and the neglect of the Stambaugh bias arising from nonstationarity in the IVX and Dlasso methods. The empirical findings echo Mankiw, (2024)’s latest pessimistic remark: “The large confidence intervals for the natural rate, together with the apparent futility of this Holy Grail search, lead me to think that we should not expect much from the Phillips curve as a guide for forecasting inflation or for judging the stance of policy.”
@@ 6 Conclusion
Mei and Shi, (2024) and the current paper respectively solve the problems of LASSO estimation and inference for predictive regressions in the presence of unit roots. In addition to this simplest form of nonstationary regressors, there are many other models that portray the complex nonstationarity in real-world data, including local unit roots, fractional integration, cointegration, structural breaks, to name a few. LASSO’s consistency and inference under these time series models are promising future research topics. Joint inference for a null of high dimensional parameters under nonstationarity is another fruitful direction.
@@ A.1 Proof of Lemma 1
$\sup_{j\in[p]}\sup_{0\leq t\leq n-1}n^{-1/2}|\xi_{jt}-\zeta_{jt}|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right).$
@@ A.2.1 Unit Root Regressors
With these preparatory lemmas, we will prove Proposition 1 stated in the main text.
Proof of Proposition 1.
@@
$\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j,\cdot}\widehat{\varphi}^{(j)}\|_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j}\varphi^{(j)*}\|_{2}^{2}+\mu\|D_{-j}\varphi^{(j)*}\|_{1}.$
for an arbitrary $(p-1)$-dimensional vector $\varphi^{(j)*}$. We can write the above inequality into
$\lVert\dfrac{W_{-j,\cdot}\left(\widehat{\varphi}^{(j)}-\varphi^{(j)*}\right)}{n}\rVert_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\leq$ $\dfrac{2}{n}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi^{(j)*})\|_{\infty}\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}+\mu\|D_{-j}\varphi^{(j)*}\|_{1}.$ (A.32)
Write the coefficient vector as $\varphi^{(j)*}=(\varphi_{k}^{(j)*})_{k\in[p],k\neq j}$, where $\varphi_{k}^{(j)*}$ is the coefficient of $w_{k,t}$. Define $\varphi_{\mathcal{M}_{x}}^{(j)*}=(\varphi_{k}^{(j)*})_{k\in\mathcal{M}_{x},k\neq j}$ and $\varphi_{\mathcal{M}_{z}}^{(j)*}=(\varphi_{z}^{(j)*})_{k\in\mathcal{M}_{z}}.$ Take a vector $\varphi^{(j)*}$ such that
$\|\varphi_{\mathcal{M}_{x}}^{(j)*}\|_{1}$ $=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+\tau\wedge(1-\tau)}}}\text{ and }\|\varphi_{\mathcal{M}_{z}}^{(j)*}\|_{1}=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.33)
for a positive number $C_{\varphi}=O(1)$. We will show that
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi^{(j)*})\|_{\infty}\leq\mu/6,$ (A.34)
@@
$\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq\dfrac{\mu}{3}\cdot\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}+\mu\|D_{-j}\varphi^{(j)*}\|_{1}\leq\dfrac{\mu}{3}\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}+\dfrac{4\mu}{3}\|D_{-j}\varphi^{(j)*}\|_{1},$
which implies $\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq 2\|D_{-j}\varphi^{(j)*}\|_{1}.$ In addition, by (A.11) and (A.12) in Lemma A.2, there exists a absolute constant $C_{\sigma}$ such that
@@
$\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$ $\leq 2\|D_{-j}\varphi^{(j)*}\|_{1}$
$\leq 2\left(\max_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{x}}^{(j)*}\|_{1}+\max_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{z}}^{(j)*}\|_{1}\right)$
$\leq\dfrac{2C_{\varphi}\cdot C_{\sigma}(\log p)^{1+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.36)
@@
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi^{(j)*}\|_{\infty}$
$\leq$ $\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{x}}^{(j)*}\|_{1}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{x}}^{(j)*}\|_{1}$
$\leq$ $\ \max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}.$
By the sup-exponential distributions of the stationary components, we can deduce that
@@
Thus, by (A.33),
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi^{(j)*}\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n(\log p)^{3}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+(\log p)^{2}\cdot\dfrac{(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$

### v2

@@ 4.1 Setup
where $R_{n}=\text{diag}(0,0.3,0.3,\ldots,0.3)$. Under this $R_{n}$ the error term $u_{t}$ in the main regression (2.1) remains i.i.d. satisfying the m.d.s. condition in Assumption 1, while the unit root generators $\Delta X_{t}$ and the stationary regressors $Z_{t}$ are AR(1) processes. The covariance matrix $\Sigma=(\Sigma_{ij})_{i,j\in[p]}$ is specified as
@@
Recall $0_{p}$ is a $p$-dimensional zero vector and $1_{p}$ is a $p$-dimensional vector of ones. The true coefficient vectors are:
@@
Both XDlasso and Dlasso involve Slasso, where the selection of tuning parameters $\lambda$ and $\mu$ affects finite sample performance. In our experiments, we employ the block 10-fold cross-validation (CV), splitting the sample into 10 equally sized chronologically ordered consecutive blocks for validations. In Theorem 1, the tuning parameters are specified as constants multiplied by the appropriate rates of convergence determined by the sample size $n$, dimensionality $p$ and the mixing condition constant $r$. As a benchmark, we also calibrate the tuning parameters following Lee et al., (2022) to examine the validity of the theoretical orders of tuning parameters specified in Theorem 1. Specifically, we perform $500$ pilot replications for each DGP, with $n_{0}=400$ , $\left(p_{x0},p_{z0}\right)=\left(100,150\right)$, and $p_{0}=p_{x0}+p_{z0}$. In each replication $q=1,2,\dots,500$, we use the 10-fold cross-validation to choose the tuning parameters $\lambda^{(q)}$ and $\mu^{(q)}$, and calibrate the constants as
$C_{{\rm m}}^{\left(q\right)}=\lambda^{(q)}n_{0}^{1/2}/(\log p_{0})^{\frac{3}{2}+\frac{1}{2r}},\ \ C_{{\rm a}}^{\left(q\right)}=\mu^{(q)}n_{0}^{[(1-\tau)\wedge\tau]/2}/(\log p_{0})^{2+\frac{1}{2r}},$
where $r=1$ and $\tau=0.5$ are chosen in the simulation. We then fix $\widehat{C}_{\star}=\mathrm{median}(C_{\star}^{\left(1\right)},\ldots,C_{\star}^{\left(500\right)})$ for $\star\in\left\{{\rm m},{\rm a}\right\}$ in the full-scale experiments. The tuning parameters are then set as $\widehat{\lambda}=\widehat{C}_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\widehat{\mu}=\widehat{C}_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ as in Theorem 1.
@@ 5 Empirical Applications
In recent years, high dimensional macroeconomic data have been extensively used to forecast key variables of interest; see Smeekes and Wijler, (2018), Medeiros et al., (2021), and Giannone et al., (2021), for example. Researchers have primarily focused on the point estimation of forecast. There has been limited empirical exploration of statistical inference on predictive power of specific predictors, due to a lack of suitable toolkits. This section showcases two empirical applications using our proposed XDlasso inference method. We utilize the monthly data of 112 U.S. macroeconomic variables spanning from January 1960 to April 2024, sourced from the FRED-MD dataset by McCracken and Ng, (2016).
@@ 5.1 Predictability of Stock Return Using Earnings-Price Ratio
In spite of a prominent economist and statistician with a national reputation, Irving Fisher suffered devastating personal finance losses during the 1929 Great Depression (Dimand and Geanakoplos,, 2005). Were he alive today, he would have mixed feelings about the fact that return predictability remains inconclusive after a century of extensively studies. Among the various predictors examined, the predictive power of valuation ratios, particularly the earnings-price ratio, has been subject to widespread discussions. In this section, we investigate the predictability of stock returns using the log earnings-price ratio while incorporating high dimensional mixed-root control covariates.
@@ 5.1.1 Data
Figure B.1 displays the time series plot of the monthly return of the S&P 500 index and the log earnings-price ratio from January 1960 to April 2024. The log earnings-price ratio exhibits persistent patterns, while the S&P 500 return appears stationary. We further report the AR(1) coefficient estimates and Augmented Dickey–Fuller (ADF) test $p$-values of both series under different sample periods in Table B.1. The S&P 500 return is evidenced to be stationary under the full sample with an ADF test $p$-value below 1%, rejecting the null hypothesis of nonstationarity. Conversely, the log earnings-price ratio shows high persistence with an AR(1) coefficient estimate equal to $0.993$. Given the $p$-value of $0.081$, nonstationarity is not rejected under the 5% significance level. Note that the nonstationary log earnings-price ratio is likely to have predictive power for the stationary monthly return of S&P 500, since we allow for a local-to-zero coefficient to balance the different scales between a stationary outcome and a nonstationary regressor (Phillips,, 2015). Theorem 2 and the paragraph that follows illuminate that inference by XDlasso has the power to detect a wide range of local-to-zero alternatives.
@@
Our analysis incorporates high dimensional covariates including all other 110 macroeconomic variables from the FRED-MD dataset. These variables are mixed with stationary and nonstationary time series. To facilitate the stationarization, FRED assigns a transformation code (TCODE) to each variable denoting the recommended transformation of the potentially nonstationary time series into a stationary one. Given the stationary pattern of the dependent variable, in our analysis, we follow the TCODE to transform the 110 time series and subsequently use the transformed variables as covariates. Nevertheless, we highlight that these elementary transformations are not a silver bullet in getting rid of nonstationarity. We perform the ADF test for each of the transformed variables for different sample periods. A nontrivial proportion (9.1%) of the transformed variables still demonstrate nonstationarity based on the ADF test at 5% significance level for both Pre-1994 and Post 1994 subperiods. The high persistence of the log earnings-price ratio and other covariates suggests the necessity of XDlasso.
@@ 5.1.2 Results
We study the one-month-ahead regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes a high dimensional vector that collects all control variables. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the log earnings-price ratio in forecasting the S&P 500 monthly return. In addition to XDlasso and Dlasso, we conduct the IVX inference in a simple regression setup, omitting all control covariates, to demonstrate the necessity of incorporating high dimensional control covariates in practice. We compare the results of XDlasso to those of Dlasso for the high dimensional regression and IVX for a simple regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+u_{t}.$
@@ 5.2 Predictability of Inflation Using Unemployment Rate
It is essential for monetary policymakers to understand the relationship between unemployment and inflation. As Engemann, (2020) pointed out, “The Federal Reserve has a dual mandate to promote maximum sustainable employment and price stability.” First alluded by Fisher, (1926, 1973) though, Phillips, (1958) popularized the Phillips curve — a negative relationship between the level of unemployment and the change rate of money wage rates. There has been a prolonged debate about whether the unemployment is a credible barometer for inflation among not only modern economic studies, but also policymakers. [^4] Mary Daly, San Francisco Fed President, delivered at Daly, (2019) a negative view on the Phillips curve that “As for the Phillips curve… most arguments today center around whether it’s dead or just gravely ill. Either way, the relationship between unemployment and inflation has become very difficult to spot.” John Williams, New York Fed President, expressed a different opinion that “The Phillips curve is the connective tissue between the Federal Reserve’s dual mandate goals of maximum employment and price stability. Despite regular declarations of its demise, the Phillips curve has endured. It is useful, both as an empirical basis for forecasting and for monetary policy analysis.” See Engemann, (2020) for more details. Empirical findings find that inflation rate can be either positively or negatively correlated to unemployment, depending on the shocks to the economies, the policies, and the lag orders (Niskanen,, 2002; Gordon,, 2011, 2013). Given the ongoing debate, We revisit the Phillips curve in a predictive regression framework utilizing the FRED-MD dataset.
@@ 5.2.1 Data
Similar to our first empirical application in Section 5.1, we follow Benati, (2015) to delineate three subperiods in addition to the full sample: Pre-Volcker (January 1960–July 1979), Volcker and Greenspan (August 1979–January 2006), and Bernanke, Yellen, and Powell (February 2006 - April 2024). These subperiods correspond to different eras in U.S. monetary policy, each named after the Federal Reserve chairperson who presided during that time. [^5] Pre-Volcker (before August 1979): This period was characterized by high and volatile inflation, with the Federal Reserve lacking a clear nominal anchor. Volcker and Greenspan (August 1979–January 2006): Volcker, who served as the Federal Reserve Chairman from August 1979 to August 1987, implemented aggressive anti-inflation measures, notably raising interest rates to historically high levels. Greenspan succeeded Volcker and continued to focus on maintaining price stability during his tenure, which contributed to a period of low and stable inflation known as the “Great Moderation”. Bernanke, Yellen, and Powell (February 2006–December 2019): Bernanke’s tenure as Fed Chairman was marked by the Great Recession and the implementation of unconventional monetary policies, such as quantitative easing, aimed at stimulating the economy and preventing deflation. Yellen and Powell carried on these policies. This periodization allows us to examine how the relationship between inflation and unemployment may have evolved across different policy regimes and economic conditions.
@@ 5.2.2 Results
We study, in our predictive regression framework, a one-month-ahead regression $\text{Inflation}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{Unrate}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes all transformed covariates. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the unemployment rate in forecasting inflation. The predictive form is of particular interest among policymakers in leveraging the relationship as a practical tool.
@@
Table 4 reports the point estimates and standard errors for $\theta_{1}^{*}$ using IVX, Dlasso and XDlasso across sample periods. The results by XDlasso in different subperiods are unanimous that the unemployment rate does not have significant predictive power in forecasting the inflation rate across all periods. These results add new insight to the recent debates on the Phillips curve: the predictive power of unemployment for inflation appears to be diminished when we control for other high dimensional time series as covariates. In contrast, IVX and Dlasso yield contradictory results. Dlasso delivers significantly positive estimates for the full sample and the Pre-Volcker period at the 1% significance level. Conversely, IVX detects significant negative effects for the full sample and Bernanke-Yellen-Powell period at the 5% and 10% significance levels, respectively. These discrepancies may be attributed to the failure of IVX to control for confounders and the neglect of the Stambaugh bias in Dlasso arising from nonstationarity. The empirical findings echo Mankiw, (2024)’s latest pessimistic remark: “The large confidence intervals for the natural rate, together with the apparent futility of this Holy Grail search, lead me to think that we should not expect much from the Phillips curve as a guide for forecasting inflation or for judging the stance of policy.”
@@ 6 Conclusion
One limitation of this paper is that it considers only the unit roots for persistent regressors. In addition to unit roots, there are many other models that portray the complex nonstationarity in real-world data, including local unit roots, fractional integration, cointegration, structural breaks, to name a few. LASSO’s consistency and inference under these time series models are yet to be studied, and one may attempt to generalize inference of the null to high dimensional parameters. These extension are currently being investigated.
@@ A.1 Proof of Lemma 1
$\sup_{j\in[p+1]}\sup_{0\leq t\leq n-1}n^{-1/2}|\xi_{jt}-\zeta_{jt}|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right).$
@@ A.2.1 Unit Root Regressors
With these preparatory lemmas, we will prove Proposition 1 for $j\in\mathcal{M}_{x}$.
Proof of Proposition 1 for $j\in\mathcal{M}_{x}$..
@@
$\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j,\cdot}\widehat{\varphi}^{(j)}\|_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j}\varphi\|_{2}^{2}+\mu\|D_{-j}\varphi\|_{1}.$
for an arbitrary $(p-1)$-dimensional vector $\varphi$. We can write the above inequality into
$n^{-1}\lVert W_{-j,\cdot}\left(\widehat{\varphi}^{(j)}-\varphi\right)\rVert_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\leq$ $\dfrac{2}{n}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi)\|_{1}+\mu\|D_{-j}\varphi\|_{1}.$ (A.32)
Write the coefficient vector as $\varphi=(\varphi_{k})_{k\in[p],k\neq j}$, where $\varphi_{k}$ is the coefficient of $w_{k,t}$. Define $\varphi_{\mathcal{M}_{x}}=(\varphi_{k})_{k\in\mathcal{M}_{x},k\neq j}$ and $\varphi_{\mathcal{M}_{z}}=(\varphi_{k})_{k\in\mathcal{M}_{z}}.$ Take a vector $\varphi$ such that
$\|\varphi_{\mathcal{M}_{x}}\|_{1}$ $=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+\tau\wedge(1-\tau)}}}\text{ and }\|\varphi_{\mathcal{M}_{z}}\|_{1}=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.33)
for a positive number $C_{\varphi}=O(1)$. Without loss of generality, suppose that $s\geq 1$. We will show that
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\leq\frac{\mu}{2}\left(1-\frac{1}{2s^{2}}\right),$ (A.34)
@@
$\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$ $\leq\mu\left[\left(1-\frac{1}{2s^{2}}\right)\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi)\|_{1}+\|D_{-j}\varphi\|_{1}\right]$
$\leq\mu\left[\left(1-\frac{1}{2s^{2}}\right)\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}+2\|D_{-j}\varphi\|_{1}\right],$
which implies $\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq 4s^{2}\|D_{-j}\varphi\|_{1}.$ In addition, by (A.11) and (A.12) in Lemma A.2, there exists a absolute constant $C_{\sigma}$ such that
@@
$\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$ $\leq 4s^{2}\|D_{-j}\varphi\|_{1}$
$\leq 4s^{2}\left(\max_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{x}}\|_{1}+\max_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{z}}\|_{1}\right)$
$\leq\dfrac{4C_{\varphi}\cdot C_{\sigma}s^{2}(\log p)^{1+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.36)
@@
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi\|_{\infty}$
$\leq$ $\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{x}}\|_{1}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{z}}\|_{1}$
$\leq$ $\ \max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}},$ (A.38)
where the second inequality applies (A.33). By the sup-exponential distributions of the stationary components, we can deduce that
@@
Thus, by (A.38), (A.40), and (A.41),
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n(\log p)^{3}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+(\log p)^{2}\cdot\dfrac{(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$